# VULCA-Bench: A Multicultural Vision-Language Benchmark for Evaluating Cultural Understanding

### v1

@@ Abstract
We introduce Vulca-Bench, a multicultural art-critique benchmark for evaluating Vision–Language Models’ (VLMs) cultural understanding beyond surface-level visual perception. Existing VLM benchmarks predominantly measure L1–L2 capabilities (object recognition, scene description, and factual question answering) while under-evaluate higher-order cultural interpretation. Vulca-Bench contains 7,410 matched image–critique pairs spanning eight cultural traditions, with Chinese–English bilingual coverage. We operationalise cultural understanding using a five-layer framework (L1–L5, from Visual Perception to Philosophical Aesthetics), instantiated as 225 culture-specific dimensions and supported by expert-written bilingual critiques. Our pilot results indicate that higher-layer reasoning (L3–L5) is consistently more challenging than visual and technical analysis (L1–L2). The dataset, evaluation scripts, and annotation tools are available under CC BY 4.0 in the supplementary materials.
@@
Haorui Yu and Ramon Ruiz-Dolz and Diji Yang and Hang He and Fengrui Zhang and Qiufeng Yi
@@ 1 Introduction
Vision-Language Models (VLMs) achieve remarkable performance on visual understanding tasks, including 90%+ on object detection Li et al. (2023) and 77.2% on VQAv2 Goyal et al. (2017). However, these benchmarks predominantly assess L1–L2 visual perception, which involves recognising objects, describing scenes, and answering factual questions. They fail to evaluate cultural understanding, namely the capacity to interpret symbolic meanings, appreciate aesthetic traditions, and engage with philosophical concepts embedded in visual content. Recent studies confirm this gap: GPT-4o achieves only 54.1% accuracy on Chinese traditional painting Zeng et al. (2024), and cross-cultural evaluations reveal significant Western bias in VLM performance Liu et al. (2025).
@@
We define cultural understanding using a five-layer framework that distinguishes levels of interpretation: L1 (visual perception), L2 (technical analysis), L3 (cultural symbolism), L4 (historical context), and L5 (philosophical aesthetics). Higher layers require progressively deeper cultural knowledge (Section 3 provides full definitions). Consider a Chinese ink painting of plum blossoms (梅花): a VLM may correctly identify “plum blossoms” and “ink wash technique” (L1–L2), yet miss the symbolic meaning of resilience against adversity (L3) Zhang (2024), the artist’s lineage within the “Four Gentlemen” tradition (L4), or the aesthetic principle of 气韵生动 (qiyun shengdong, “spirit resonance”) that defines Chinese painting philosophy (L5) Jing (2023). This example illustrates that cultural understanding is not a single capability but a progression across interpretive levels, a progression that existing benchmarks fail to measure.
@@
Recent cultural AI datasets (CulturalBench Chiu et al. (2025), CultureAtlas Fung et al. (2024), GIMMICK Schneider et al. (2025)) begin to address this gap. Initial work on Chinese painting critique Yu et al. (2025a) introduced quantitative VLM evaluation using 163 expert commentaries, revealing significant VLM-expert divergence. Complementary probing Yu et al. (2025b) found VLMs rely on “symbolic shortcuts,” excelling on Western festivals but failing on underrepresented traditions. However, these studies focused on single cultures; systematic cross-cultural evaluation remains absent. Existing datasets remain L1-centric (over 90% of metrics focus on visual perception), lack hierarchical frameworks distinguishing basic visual analysis from higher-order symbolic reasoning, and exhibit Western bias by underrepresenting Asian, Middle Eastern, and South Asian traditions.
@@
We introduce Vulca-Bench (Vision-Understanding-Language-Culture Assessment), a comprehensive multicultural art critique benchmark addressing these gaps. Vulca-Bench provides 7,410 matched image-critique pairs across 8 traditions with 100% bilingual coverage and 98% cultural-fact accuracy, operationalised through our five-layer framework (L1–L5) spanning 225 culture-specific dimensions that preserve Cultural Symmetry (i.e., equal methodological treatment across cultures regardless of sample size). All critiques meet expert-quality standards (Chinese $\geq$150 characters, English $\geq$100 words) with cultural specialist validation. The complete dataset is available under CC BY 4.0 in the supplementary materials.
Vulca-Bench enables systematic evaluation of VLMs’ hierarchical cultural understanding, identification of L1–L5 performance gaps, equal-weighted cultural fairness probing on the balanced-pilot subset (N=336, 7 cultures), and development of culturally grounded VLM architectures. Figure 1 showcases representative artworks from each tradition.
@@
This paper makes three contributions. (1) We present Vulca-Bench, a multicultural art-critique benchmark enabling evaluation of cultural interpretation beyond surface-level perception, comprising 7,410 image–critique pairs across eight traditions with 225 expert-defined dimensions. (2) We formalise the Cultural Symmetry Principle, which enforces schema and protocol parity across cultures (without requiring equal sample sizes), thereby supporting fair cross-cultural evaluation. (3) We provide pilot experiments and error analyses showing that Vulca-Bench exposes systematic failures in higher-layer cultural reasoning (L3–L5) that are not captured by standard VLM benchmarks.
@@ 2 Related Work
Standard VLM benchmarks focus on L1 perception: MME Fu et al. (2023) and SEED-Bench-2 Li et al. (2024) achieve 90%+ on object detection but lack cultural evaluation; POPE Li et al. (2023) tests object existence without symbolic significance; and specialised benchmarks such as MathVista Lu et al. (2024), ChartQA Masry et al. (2022), and DocVQA Mathew et al. (2021) evaluate domain reasoning but not cultural understanding. Cultural AI datasets have begun to address this gap, yet remain limited. CulturalBench Chiu et al. (2025) and CulturalVQA Nayak et al. (2024) adopt QA formats rather than generative critique; GIMMICK Schneider et al. (2025) reveals Western biases across 144 countries but employs recognition tasks; WuMKG Wan et al. (2024) provides a Chinese painting knowledge graph (104K entities) for retrieval, not generative evaluation; and CulTi Yuan et al. (2025) offers image-text pairs for heritage retrieval without hierarchical cultural dimensions.
Art and aesthetics datasets similarly fall short. WikiArt Saleh and Elgammal (2015) (77K paintings, 27 styles) and OmniArt Strezoski and Worring (2018) (429K artworks) enable style classification but lack expert critiques or cultural dimension annotations; ArtEmis Achlioptas et al. (2021) focuses on affective responses; and Zhang et al. Zhang (2024) model Chinese painting critique via Xie He’s Six Principles at L1–L2 only. Prior VLM art critique work Yu et al. (2025a, b) focused on single cultures or specific visual domains, leaving critical gaps: no hierarchical L1–L5 schema operationalised across multiple cultures, no non-Western aesthetic frameworks such as Chinese yijing or Japanese wabi-sabi, and no expert-quality generative critique data with cultural dimension coverage.
Vulca-Bench addresses all three gaps through cross-cultural extension to eight traditions. Unlike QA-based cultural benchmarks that test factual recall, or recognition tasks that assess surface-level identification, Vulca-Bench evaluates generative cultural critique, requiring VLMs to produce coherent interpretations spanning visual perception (L1–L2) through philosophical aesthetics (L5). Our 225 culture-specific dimensions enable fine-grained diagnostic evaluation (layer-wise drop analysis, dimension coverage scoring) unavailable in existing datasets.
@@ 3.1 Cultural Symmetry Principle
Because Western and Chinese traditions constitute the majority of the full corpus (82%), we release two balanced variants for fair comparison: Balanced (N=384; 48 per culture $\times$ 8 cultures) and Balanced-Pilot (N=336; 48 per culture $\times$ 7 cultures, excluding Hermitage). We recommend using balanced variants for per-culture comparisons and the full corpus for aggregate benchmarking with tighter confidence intervals.
@@
The 70% threshold derives from pilot annotator calibration: critiques covering fewer than 70% of dimensions typically lacked substantive L3–L5 engagement. We therefore set this threshold empirically, based on the observation that lower coverage correlated strongly with missing philosophical or historical analysis. Pairs falling below this threshold are either revised by annotators or excluded from the corpus (127 pairs removed, 1.8% of candidates, distributed proportionally across cultures).
Beyond structural parity, we validate measurement alignment using three criteria. First, layer-wise difficulty exhibits a consistent monotonic decline from L1 to L5 across cultures; the resulting five-point layer profiles are highly similar (mean pairwise Pearson correlation across cultures $r=0.96$). Second, inter-annotator agreement is comparable across cultures: Cohen’s $\kappa=0.79$ for L1–L2 and $0.71$ for L3–L5, with cross-culture standard deviation below 0.04. Third, each culture’s L5 captures its deepest interpretive challenge, supporting functional equivalence. See Appendix B for a finer-grained description of the corpus (i.e., complete, evaluation, gold, human, and balanced splits).
@@ 3.2 Cultural Understanding Framework
We operationalise cultural understanding through our five-layer framework reflecting increasing interpretive complexity. L1–L2 require visual observation; L3–L5 demand progressively deeper cultural knowledge that VLMs typically lack.
@@
Each culture adapts L1–L5 to its indigenous aesthetic vocabulary, yielding 225 dimensions total. Dimension counts (25–30) reflect tradition complexity: Chinese includes 30 dimensions due to rich philosophical aesthetics (e.g., qiyun 气韵, yijing 意境); Western 25 dimensions emphasise formal analysis (chiaroscuro, perspective). All critiques must cover $\geq$70% of culture-specific dimensions, ensuring substantive L3–L5 engagement.
@@
Western 25 4,041 Chiaroscuro, linear perspective, impasto
@@
Japanese 27 401 Wabi-sabi, yugen, ma (negative space)
Islamic 28 205 Geometric patterns, arabesques, Persian miniature
Mural 30 201 Buddhist art, Dunhuang murals, cave painting
Hermitage 30 196 Russian/European court art, iconography
Indian 30 173 Rasa theory (9 emotions), shringara, bhakti
Korean 25 151 Literati aesthetics, muninhwa tradition
@@ 3.3 Data Sources
We curate artworks from authoritative museum collections with open-access policies. Chinese artworks derive from the Palace Museum, Shanghai Museum, and Metropolitan Museum; Western works from the Metropolitan Museum, Louvre, and Hermitage Museum; Japanese from Tokyo and Kyoto National Museums; Korean from the National Museum of Korea; Islamic from the Museum of Islamic Art Doha and Metropolitan Museum; Indian from the National Museum Delhi and British Museum; and Hermitage works from the State Hermitage Museum collection (Russian and European court art, religious iconography). All images are released under CC BY 4.0 in the supplementary materials.
@@ 3.4 Bilingual Critique Annotation
We adopt Chinese–English bilingual critiques for three reasons. First, terminology preservation: many Chinese aesthetic concepts (e.g., qiyun 气韵 “spirit resonance,” yijing 意境 “artistic conception”) are more precisely expressed in Chinese and are critical for consistent dimension annotation and automated matching. Second, international accessibility: English enables broader scholarly use while retaining romanised culture-specific terms where appropriate. Third, expert availability: our annotation team (Table 3) comprises bilingual specialists with deep cultural expertise, enabling controlled drafting, translation, and review within a single protocol. This design reflects practical constraints rather than linguistic preference; extending to additional native-language critiques (e.g., Japanese, Korean, Arabic, Hindi/Sanskrit) is a natural next step discussed in Limitations.
@@
Expert annotators generate critiques following the L1–L5 framework. Table 3 summarises the annotation team. Each annotator receives the artwork image and metadata (title, artist, period), then generates a critique covering all 5 layers with culture-specific dimensions. Chinese critiques must reach $\geq$150 characters (average 450), and English translations preserve cultural terminology with romanisation. A reviewer validates each critique on dimension coverage ($\geq$70%), cultural accuracy, and bilingual consistency. Annotator-reviewer disagreements are resolved through discussion with reference to authoritative sources; unresolved cases (<<2%) escalate to senior expert panel. Full annotation guidelines appear in Appendix E; representative bilingual critiques demonstrating L1–L5 coverage are provided in Appendix F.
@@
Korean follows a modified protocol due to the scarcity of qualified Korean art-history specialists with dual fluency. Rather than compromise annotation quality, we employ one primary annotator (PhD in Korean studies; 15+ years of specialisation in Joseon-era painting) with 100% expert review by a senior scholar. The reported $\kappa$ therefore measures annotator–reviewer agreement rather than independent dual-annotation IAA; we report Korean separately for methodological transparency.
@@
Table 4: Inter-annotator agreement by culture and layer. L3–L5 shows lower but substantial agreement, reflecting interpretive complexity. Pooled excludes Korean (5 cultures, N=90). Korean uses a different protocol: 1 primary annotator (PhD Korean studies) + 100% expert review. The reported $\kappa$ is annotator–reviewer agreement, not independent dual-annotation IAA; we report it separately for quality transparency but do not include it in pooled statistics.
3.5 Quality Assurance
Multi-stage validation ensures data quality. Each record contains an explicit dimension checklist through its covered_dimensions field, listing dimension IDs (e.g., ["CN_L1_D1", "CN_L3_D2", ...]) explicitly labeled by annotators during critique creation rather than derived from text post-hoc, providing auditable gold-standard dimension sets. Automated checks enforce the $\geq$70% dimension coverage threshold (21/30 for CN, 18/25 for WE, etc.) based on covered_dimensions length. We removed all exact duplicates after deep audit. Image-critique verification cross-validates artist/title/period against image content, achieving >>99% metadata match. All pairs have both Chinese ($\geq$150 characters) and English critiques, ensuring bilingual completeness. Finally, all images are compressed to $\leq$3.75MB for VLM API compatibility.

### v2

@@ Abstract
We introduce Vulca-Bench, a multicultural art-critique benchmark for evaluating Vision–Language Models’ (VLMs) cultural understanding beyond surface-level visual perception. Existing VLM benchmarks predominantly measure L1–L2 capabilities (object recognition, scene description, and factual question answering) while under-evaluate higher-order cultural interpretation. Vulca-Bench contains 7,410 matched image–critique pairs spanning eight cultural traditions, with Chinese–English bilingual coverage. We operationalise cultural understanding using a five-layer framework (L1–L5, from Visual Perception to Philosophical Aesthetics), instantiated as 225 culture-specific dimensions and supported by expert-written bilingual critiques. Our pilot results indicate that higher-layer reasoning (L3–L5) is consistently more challenging than visual and technical analysis (L1–L2). The dataset, evaluation scripts, and annotation tools are available under CC BY 4.0 on GitHub [^1] https://github.com/yha9806/VULCA-Bench and Hugging Face [^2] https://huggingface.co/datasets/harryHURRY/vulca-bench.
@@
Haorui Yu1, Ramon Ruiz-Dolz2, and Diji Yang3, and Hang He4, and Fengrui Zhang5, and Qiufeng Yi6 1DJCAD, University of Dundee, United Kingdom 2Faculty of Science, Engineering and Business, University of Dundee, United Kingdom 3 University of California, Santa Cruz, USA 4East China Normal University, China 5Nanjing University, China 6School of Computer Science, University of Birmingham, United Kingdom
@@ 1 Introduction
Vision-Language Models (VLMs) achieve remarkable performance on visual understanding tasks. For instance, on object hallucination benchmarks like POPE Li et al. (2023) where leading models such as InstructBLIP reach 88.7% accuracy, and on visual question answering such as VQAv2 Goyal et al. (2017) where best models exceed 75% accuracy. However, these benchmarks predominantly assess visual perception, which involves recognising objects, describing scenes, and answering factual questions. They fail to evaluate cultural understanding, namely the capacity to interpret symbolic meanings, appreciate aesthetic traditions, and engage with philosophical concepts embedded in visual content. Recent studies confirm this gap: GPT-4o achieves only 54.1% overall accuracy on CII-Bench, with even lower performance (51.8%) on Chinese Traditional Culture content Zhang et al. (2025), and cross-cultural evaluations reveal significant Western bias in VLM performance Liu et al. (2025).
@@
We define cultural understanding by using a five-layer framework that distinguishes levels of interpretation: L1 (visual perception), L2 (technical analysis), L3 (cultural symbolism), L4 (historical context), and L5 (philosophical aesthetics). Higher layers require progressively deeper cultural knowledge (see Section 3 for complete definitions). Consider a Chinese ink painting of plum blossoms (梅花): a VLM may correctly identify “plum blossoms” and “ink wash technique” (L1–L2), yet miss the symbolic meaning of resilience against adversity (L3), the artist’s lineage within the “Four Gentlemen” (四君子) tradition (L4), or aesthetic principles such as 气韵生动 (qiyun shengdong, “spirit resonance”) and 意境 (yijing, “artistic conception”) Jing (2023) that define Chinese painting philosophy (L5). This example illustrates that cultural understanding is not a single capability but a progression across interpretive levels, a progression that existing benchmarks fail to measure.
@@
Recent cultural AI datasets (e.g., CulturalBench Chiu et al. (2025), CultureAtlas Fung et al. (2024), GIMMICK Schneider et al. (2025)) begin to address this gap. Initial work on Chinese painting critique Yu et al. (2025a) introduced a quantitative VLM evaluation method using 163 expert commentaries, revealing significant VLM-expert divergence. Complementary probing Yu et al. (2025b) found VLMs rely on “symbolic shortcuts”, excelling on Western festivals but failing on under-represented traditions. However, these studies focused on single cultures; systematic cross-cultural evaluation remains absent. Furthermore, existing datasets remain L1-centric (over 90% of metrics focus on visual perception), lack hierarchical frameworks distinguishing basic visual analysis from higher-order symbolic reasoning, and exhibit Western bias by underrepresenting Asian, Middle Eastern, and South Asian traditions.
@@
This paper introduces Vulca-Bench (Vision-Understanding and Language-based Cultural Assessment), a comprehensive multicultural art critique benchmark addressing these gaps. Vulca-Bench provides 7,410 matched image-critique pairs across 8 traditions with 100% bilingual coverage and expert-validated factual accuracy, annotated using our five-layer framework (L1–L5) spanning 225 culture-specific dimensions that preserve Cultural Symmetry (i.e., equal methodological treatment across cultures regardless of sample size). All critiques meet expert-quality standards (Chinese $\geq$150 characters, English $\geq$100 words) with cultural specialist validation. Vulca-Bench enables systematic evaluation of VLMs’ hierarchical cultural understanding, identification of L1–L5 performance gaps, equal-weighted cultural fairness probing on the balanced-pilot subset (N=336, 7 cultures), and development of culturally grounded VLM architectures. Figure 1 showcases representative artworks from each tradition.
@@
This paper therefore makes three contributions. (1) We present Vulca-Bench, a multicultural art-critique benchmark enabling evaluation of cultural interpretation beyond surface-level perception, comprising 7,410 image–critique pairs across eight traditions with 225 expert-defined dimensions. (2) We formalise the Cultural Symmetry Principle, which enforces schema and protocol parity across cultures (without requiring equal sample sizes), thereby supporting fair cross-cultural evaluation. (3) We provide pilot experiments and error analyses showing that Vulca-Bench exposes systematic failures in higher-layer cultural reasoning (L3–L5) that are not captured by standard VLM benchmarks.
@@ 2 Related Work
Standard VLM benchmarks focus on L1 perception: MME Fu et al. (2025) evaluates 14 subtasks spanning perception and cognition, SEED-Bench-2 Li et al. (2024) covers 27 dimensions with 24K multiple-choice questions, and POPE Li et al. (2023) probes object hallucination through binary yes/no questions. However, none address cultural understanding in-depth. Specialised benchmarks such as MathVista Lu et al. (2024), ChartQA Masry et al. (2022), and DocVQA Mathew et al. (2021) evaluate domain reasoning but not cultural interpretation. Cultural AI datasets have begun to address this gap, yet remain limited. CulturalBench Chiu et al. (2025) and CulturalVQA Nayak et al. (2024) adopt QA formats rather than generative critique; GIMMICK Schneider et al. (2025) reveals Western biases across 144 countries employing recognition tasks; WuMKG Wan et al. (2024) provides a Chinese painting knowledge graph (104K entities) useful for information retrieval rather than generative evaluation; and CulTi Yuan et al. (2025) offers image-text pairs for heritage retrieval without hierarchical cultural dimensions.
Art and aesthetics datasets similarly fall short. WikiArt Saleh and Elgammal (2016) ($\sim$80K paintings, 27 styles) and OmniArt Strezoski and Worring (2018) (>>1M artworks) enable style classification but lack expert critiques or cultural dimension annotations; ArtEmis Achlioptas et al. (2021) focuses on affective responses; and Zhang et al. (2024) survey computational approaches for traditional Chinese painting based on Xie He’s Six Principles, focusing primarily on visual and technical analysis. Prior VLM art critique work Yu et al. (2025a, b) focused on single cultures or specific visual domains, leaving critical gaps: no hierarchical L1–L5 schema operationalised across multiple cultures, no non-Western aesthetic frameworks such as Chinese yijing or Japanese wabi-sabi, and no expert-quality generative critique data with cultural dimension coverage.
Vulca-Bench addresses all three gaps through cross-cultural extension to eight traditions. Unlike QA-based cultural benchmarks that test factual recall or recognition tasks that assess surface-level identification, Vulca-Bench evaluates generative cultural critique, requiring VLMs to produce coherent interpretations spanning visual perception (L1–L2) through philosophical aesthetics (L5). Our 225 culture-specific dimensions enable fine-grained diagnostic evaluation (e.g., layer-wise analysis or dimension coverage scoring) unavailable in existing datasets.
@@ 3.1 Cultural Symmetry Principle
Because Western and Chinese traditions constitute the majority of the full corpus (82%), we release two balanced variants for fair comparison: Balanced (N=384; 48 per culture $\times$ 8 cultures) and Balanced-Pilot (N=336; 48 per culture $\times$ 7 cultures, excluding Hermitage, whose European court art overlaps with Western). We recommend using balanced variants for cross-culture comparisons and the full corpus for aggregate benchmarking with tighter confidence intervals.
@@
The 70% threshold derives from pilot annotator calibration: critiques covering fewer than 70% of dimensions typically lacked substantive L3–L5 engagement. We therefore set this threshold empirically, based on the observation that lower coverage correlated strongly with missing philosophical or historical analysis. Pairs falling below this threshold are either revised by annotators or excluded from the corpus (127 pairs removed, 1.8% of candidates distributed proportionally across cultures).
Beyond structural parity, we validate measurement alignment using three criteria. First, layer-wise difficulty exhibits a consistent monotonic decline from L1 to L5 across cultures. All cultures show the same pattern: L1 easiest, L5 hardest ($r=0.96$ cross-culture correlation). Second, inter-annotator agreement is comparable across cultures: Cohen’s $\kappa=0.79$ for L1–L2 and $0.71$ for L3–L5, with cross-culture standard deviation below 0.04. Third, each culture’s L5 captures its deepest interpretive challenge, supporting functional equivalence. See Appendix B for a finer-grained description of the corpus (i.e., complete, evaluation, gold, human, and balanced splits).
@@ 3.2 Cultural Understanding Framework
We operationalise cultural understanding through our five-layer framework reflecting increasing interpretive complexity. While L1–L2 require visual observation, L3–L5 demand progressively deeper cultural and contextual knowledge.
@@
Each culture adapts L1–L5 to its indigenous aesthetic vocabulary, yielding 225 dimensions total. Dimension counts (25–30) reflect tradition complexity. For example, Chinese includes 30 dimensions due to rich philosophical aesthetics (e.g., qiyun 气韵, yijing 意境). On the other hand, Western cultures cover 25 dimensions that emphasise more formal analysis (chiaroscuro, perspective). In Vulca-Bench, all critiques must cover at least 70% of their respective culture-specific dimensions, ensuring substantive L3–L5 engagement.
@@
Western 25 4,041 Chiaroscuro, perspective, impasto
@@
Japanese 27 401 Wabi-sabi, yugen, ma
Islamic 28 205 Geometric patterns, arabesques
Mural 30 201 Buddhist art, Dunhuang murals
Hermitage 30 196 Russian/European court art
Indian 30 173 Rasa theory, shringara, bhakti
Korean 25 151 Literati aesthetics, muninhwa
@@ 3.3 Data Sources
We curate artworks from authoritative museum collections with open-access policies. Chinese artworks derive from the Palace Museum, Shanghai Museum, and Metropolitan Museum; Western works from the Metropolitan Museum, Louvre, and Hermitage Museum; Japanese from Tokyo and Kyoto National Museums; Korean from the National Museum of Korea; Islamic from the Museum of Islamic Art Doha and Metropolitan Museum; Indian from the National Museum Delhi and British Museum; and Hermitage works from the State Hermitage Museum collection (Russian and European court art, religious iconography). All images are released under CC BY 4.0 at https://github.com/yha9806/VULCA-Bench and HuggingFace: https://huggingface.co/datasets/harryHURRY/vulca-bench.
@@ 3.4 Bilingual Critique Annotation
We adopt Chinese–English bilingual critiques for three reasons. First, terminology preservation: many Chinese aesthetic concepts (e.g., qiyun 气韵 “spirit resonance,” yijing 意境 “artistic conception”) are more precisely expressed in Chinese and are critical for consistent dimension annotation and automated matching. Second, international accessibility: English enables broader scholarly use while retaining romanised culture-specific terms where appropriate. Third, expert availability: our annotation team (Table 3) comprises bilingual specialists with deep cultural expertise, enabling controlled drafting, translation, and review within a single protocol. This design reflects practical constraints rather than linguistic preference, thus extending to additional native-language critiques (e.g., Japanese, Korean, Arabic, Hindi/Sanskrit) is a natural next step discussed in the Limitations section.
@@
Expert annotators generate critiques following the L1–L5 framework. Table 3 summarises the annotation team. Each annotator receives the artwork image and metadata (title, artist, period), then generates a critique covering all 5 layers with culture-specific dimensions. Chinese critiques must reach $\geq$150 characters (average 450), and English translations preserve cultural terminology with romanisation. A reviewer validates each critique on dimension coverage ($\geq$70%), cultural accuracy, and bilingual consistency. Annotator-reviewer disagreements are resolved through discussion with reference to authoritative sources. The complete scheme of our annotation guidelines can be found in Appendix E and representative bilingual critiques demonstrating L1–L5 coverage are provided in Appendix F.
@@
Korean follows a modified protocol due to the scarcity of qualified Korean art-history specialists with dual fluency. Rather than compromise annotation quality, we employ one primary annotator (PhD in Korean studies; 15+ years of specialisation in Joseon-era painting) with a 100% expert review by a senior scholar. The reported $\kappa$ therefore measures annotator–reviewer agreement rather than independent dual-annotation IAA.
@@
Table 4: Inter-annotator agreement by culture and layer. L3–L5 shows lower but substantial agreement, reflecting interpretive complexity. Pooled excludes Korean (5 cultures, N=90). Korean uses a different protocol: 1 primary annotator (PhD Korean studies) + 100% expert review.